Double Op Amps: solve for G=Vo/Vin
22 = 362 MA 27-342 M Vin = 13V 9 3R = 540 w R=182 in R=18.02 it 1918 0.12 R = 18.12 ↑ iz = lov LOV = V2 o 18.12.
Answer #1

70 R in va 3R vd 2R Vb = 13 V R + vo Vc 2 lov - V> R R=182 Vazo. &R=362 3R=542 sah:- a The op-amp-1 in inverting type op ampev ( 13 12 + 2R Vb. 3R + vo ZR. ud(a 2R+ 3R ARV + 3RVO GR ( 2% + 36) = 5R 2 b + 2 o VE 2 Vb + 3 Vo -u vint 3 Vo 1 lo 5 For op-6.5 Vin + 2.582 Vo= 3 as vo 6.5 Vin + 2.5 V2 - Vin 3 Vin 6.5t 2.5 V Vin 3 6.5 4 2.5*( 3 = 2.807692 vo 2.807692. i gain Go Jin